Croy Inc. has the following projected sales for the next five months:

Month Sales in Units

April 3,470

May 3,890

June 4,520

July 4,170

August 3,980

Croy’s finished goods inventory policy is to have 70 percent of the next month’s sales on hand at the end of each month. Direct material costs $3.00 per pound, and each unit requires 2 pounds. Raw materials inventory policy is to have 50 percent of the next month’s production needs on hand at the end of each month. Raw materials on hand at March 31 totaled 3,764 pounds. Required:

1. Determine budgeted production for April, May, and June.

2. Determine the budgeted cost of materials purchased for April, May, and June.
Answer #1

1.

Budgeted Production
April May June
Budgeted sales                              3,470                              3,890                              4,520
Add: Desired ending inventory 3,890*70% = 2,723 4,520*70% = 3,164 4,170*70% = 2,919
Total units needed                              6,193                              7,054                              7,439
Less: Budgeted begining inventory 3,470*70% = 2,429 3,890*70% = 2,723 4,520*70% = 3,164
Units to be produced                              3,764                              4,331                              4,275

2.

Budgeted cost of material to be purchase
April May June
units to be produced                              3,764                              4,331                                                                                      4,275
Raw material required per unit                                      2                                      2                                                                                              2
Total raw material required                              7,528                              8,662                                                                                      8,550
Add: Ending raw material inventory 8,662*50% = 4,331 8,550*50% = 4,275 (((4,170+(3,980*70%)-2,919)*2)*50%) = 4,037
Raw material needed                            11,859                            12,937                                                                                    12,587
Less: Beginning raw material inventory                              3,764 8,662*50% = 4,331 8,550*50% = 4,275
Raw material to be purchase                              8,095                              8,606                                                                                      8,312
Material price per pound $                                  3 $                                  3 $                                                                                          3
Total raw material cost $                       24,285 $                       25,818 $                                                                               24,936